at noon a train leaves washington d.c. headed for charleston south carolina, a distance of 500 mi. the train travels at a speed of 60mph.at 1 pm a second train leaves charleston headed for washington dc traveling at 50mph . how long after the train leaves charleston will the 2 trains pass each other
Found 2 solutions by stanbon, calegebra:
Answer by stanbon(75887)

at noon a train leaves washington d.c. headed for charleston south carolina, a distance of 500 mi. the train travels at a speed of 60mph.
---------------------------------------
at 1 pm a second train leaves charleston headed for washington dc traveling at 50mph .
---------------------------------------
how long after the train leaves charleston will the 2 trains pass each other
---
Wash to Char train DATA:
rate = 60 mph: time = x hrs; distance = rt = 60x miles
-----------------
Char to Wash train DATA:
rate = 50 mph; time = x-1 hrs ; distance = rt = 50(x-1) miles
---------------------
Equation:
dist + dist = 500 miles
60x + 50x -50 = 500
110x = 550
x = 5 hrs
-----
x-1 = 4 hrs (time for the Char to Wash train
=======================
Cheers,
Stan H.

Answer by calegebra(14)

Solve by chart method
Distance Travelled
Time Train 60mph Train 50 mph
Noon 0 0
1PM 60 0
2PM 120 50
3PM 180 100
4PM 240 150
5PM 300 200
6PM 360 250
7PM 420 300
8PM 480 350
9PM 540 400
10PM 600 450

Answer is when both columns total 500 miles so, answer is 5PM